# Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents

### v1

@@ 05-09 Mining and extraction 59 31 9
10-32 Manifacture 9,602 2,110 426
33 Repair and installation of machinery and equipment 169 39 4
35 Electricity, gas, steam and air conditioning supply 96 70 19
@@ 36-39 Water, sewerage and waste management 100 69 19
41-43 Construction, civil engineering and special activities 388 118 17
45-47 Wholesale and retail trade 1,520 275 50
49-51 Transport 35 10 1
52 Warehousing and support activities for transportation 40 9 0
53 Postal and courier activities 6 2 1
55 Accommodation 3 0 0
56 Food and beverage service activities 12 2 0
58 Publishing activities 67 3 1
59-63 ICT, motion pictures, television and broadcasting 383 51 7
64-68 Financial, insurance and real estate activities 391 87 17
69 Legal and accounting activities 21 8 3
70 Activities of head offices; management and consultancy 434 141 33
71 Architectural and engineering activities 636 207 54
72 Scientific research and development 874 300 84
73 Advertising and market research 24 1 0
74 Other professional, scientific and technical activities 115 34 8
75 Veterinary activities 2 0 0
77 Rental and leasing activities 78 12 6
78 Employment activities 7 2 2
79 Travel agency and tour operator 3 1 1
80 Security and investigation activities 17 3 0
81 Services to buildings and landscape activities 17 4 0
82 Office administrative and office support 76 22 3
@@ 84 Public administration and defence 4 1 0
85 Education 92 56 26
86 Human health activities 82 12 0
87 Residential care activities 4 0 0
88 Social work activities without accommodation 2 1 0
90-91 Cultural and creative activities 3 0 0
92-93 Sports, amusement and gambling activities 12 1 1
94 Activities of membership organisations 7 4 0
95 Repair of computers and personal goods 8 0 0
96 Other personal service activities 17 2 1
@@ B.2. Classification results
A01 Agriculture;forestry; animal husbandry; hunting;trapping; fishing 917,961 246,215 49,224 19.99
A21 Baking; edible doughs 46,654 1,742 383 21.99
A22 Butchering; meat treatment; processing poultry or fish 26,671 1,719 213 12.39
A23 Foods or foodstuffs; treatment thereof, not covered by other classes 340,402 35,332 6,004 16.99
A24 Tobacco; cigars; cigarettes; smokers’ requisites 67,251 3,230 399 12.35
A41 Wearing apparel 150,232 2,533 510 20.13
A42 Headwear 25,377 604 129 21.36
A43 Footwear 74,236 1,211 154 12.72
A44 Haberdashery; jewellery 48,611 863 76 8.81
A45 Hand or travelling articles 172,126 5,334 1,525 28.59
A46 Brushware 32,592 478 75 15.69
A47 Furniture;domestic articles or appliances; coffee mills; spice mills;suction cleaners in general 962,654 34,013 5,857 17.22
A61 Medical or veterinary science; hygiene 1,953,254 99,124 12,838 12.95
A62 Life-saving; fire-fighting 116,182 12,315 2,151 17.47
A63 Sports; games; amusements 376,895 11,390 896 7.87
@@
B01 Physical or chemical processes or apparatus in general 1,332,341 610,168 115,244 18.89
B02 Crushing, pulverising, or disintegrating; preparatory treatment of grain for milling 228,865 36,524 7,200 19.71
B03 Separation of solid materials using liquids or using pneumatic tables or jigs 68,908 29,003 5,716 19.71
B04 Centrifugal apparatus or machines for carrying-out physical or chemical processes 28,460 4,359 590 13.54
B05 Spraying or atomising in general; applying liquids or other fluent materials to surfaces, in general 284,963 31,853 4,729 14.85
B06 Generating or transmitting mechanical vibrations in general 10,698 398 42 10.55
B07 Separating solids from solids; sorting 198,511 18,846 3,011 15.98
B08 Cleaning 437,682 71,352 14,924 20.92
B09 Disposal of solid waste; reclamation of contaminated soil 57,347 55,800 14,642 26.24
B21 Mechanical metal-working without essentially removing material; punching metal 426,356 16,889 1,305 7.73
B22 Casting; powder metallurgy 191,769 23,468 1,587 6.76
B23 Machine tools; metal-working not otherwise provided for 934,873 47,275 3,842 8.13
B24 Grinding; polishing 272,096 14,666 1,182 8.06
B25 Hand tools; portable power-driven tools; tool handles 435,111 15,778 1,193 7.56
B26 Hand cutting tools; cutting; severing 198,219 10,057 1,337 13.29
B27 Working or preserving wood or similar material; nailing or stapling machines in general 83,071 3,920 528 13.47
B28 Working cement, clay, or stone 171,772 12,834 2,064 16.08
B29 Working of plastics; working of substances in a plastic state in general 541,241 52,185 9,481 18.17
B30 Presses 67,789 11,141 2,094 18.8
B31 Making paper articles; working paper 54,608 1,711 249 14.55
B32 Layered products 271,499 29,370 8,669 29.52
B33 Additive manufacturing technology 57,978 14,547 637 4.38
B41 Printing; lining machines; typewriters; stamps 212,671 7,487 1,182 15.79
B42 Bookbinding; albums; files; special printed matter 47,294 870 61 7.01
B43 Writing or drawing implements; bureau accessories 69,636 615 95 15.45
B44 Decorative arts 52,718 1,838 337 18.34
B60 Vehicles in general 952,473 189,175 50,907 26.91
B61 Railways 93,654 9,136 1,133 12.4
B62 Land vehicles for travelling otherwise than on rails 396,455 27,170 4,700 17.3
B63 Ships or other waterborne vessels; related equipment 129,617 32,229 6,311 19.58
@@
B64 Aircraft; aviation; cosmonautics 151,156 23,886 3,183 13.33
B65 Conveying; packing; storing; handling thin or filamentary material 1,509,769 135,761 19,708 14.52
B66 Hoisting; lifting; hauling 340,274 13,221 1,422 10.76
B67 Opening or closing bottles, jars or similar containers; liquid handling 65,328 2,483 231 9.3
B68 Saddlery; upholstery 4,395 81 9 11.11
B81 Microstructural devices or systems, e.g. micromechanical devices 18,441 945 58 6.14
B82 Nanotechnology 71,511 19,307 3,251 16.84
B99 Subject matter not otherwise provided for in section b 5 1 0 0
C01 Inorganic chemistry 186,032 76,230 16,775 22.01
C02 Treatment of water, waste water, sewage, or sludge 424,551 424,548 154,477 36.39
C03 Glass; mineral or slag wool 84,477 15,830 2,301 14.54
C04 Cements; concrete; artificial stone; ceramics; refractories 123,798 37,915 9,124 24.06
C05 Fertilisers; manufacture thereof 43,951 30,732 7,759 25.25
C06 Explosive; matches 7,239 405 41 10.12
C07 Organic chemistry 393,922 99,795 14,075 14.1
C08 Organic macromolecular compounds; their preparation or chemical working-up 383,504 54,643 11,400 20.86
C09 Dyes; paints, polishes, natural resins, adhesives; compositions not otherwise provided for 288,284 49,263 7,700 15.63
C10 Petroleum, gas or coke industries, technical gases containing carbon monoxide; fuels; lubricant; peat 121,224 50,492 16,968 33.61
C11 Animal or vegetable oils, fats, fatty substances or waxes; fatty acids therefrom; detergents; candles 49,325 8,741 2,350 26.88
C12 Biochemistry; beer; spirits; wine; vinegar; microbiology; enzymology; mutation or genetic engineering 369,038 85,329 17,249 20.21
C13 Sugar industry 4,211 942 498 52.87
C14 Skins; hides; pelts; leather 9,731 422 65 15.4
C21 Metallurgy of iron 120,894 28,226 3,912 13.86
C22 Metallurgy; ferrous or non-ferrous alloys; treatment of alloys or non-ferrous metals 143,994 43,115 7,228 16.76
C23 Coating metallic material; coating material with metallic material; chemical surface treatment 159,252 21,942 3,036 13.84
C25 Electrolytic or electrophoretic processes; apparatus therefor 92,911 25,903 4,850 18.72
C30 Crystal growth 31,655 11,996 1,825 15.21
C40 Combinatorial chemistry; libraries, e.g. chemical libraries, dna libraries 4,867 570 43 7.54
@@
D01 Natural or man-made threads or fibres; spinning 84,316 6,579 1,104 16.78
D02 Yarns; mechanical finishing of yarns or ropes; warping or beaming 26,954 1,038 134 12.91
D03 Weaving 40,316 1,175 162 13.79
D04 Braiding; lace-making; knitting; trimmings; non-woven fabrics 58,298 3,603 530 14.71
D05 Sewing; embroidering; tufting 42,532 548 30 5.47
D06 Treatment of textiles or the like; laundering; flexible materials not otherwise provided for 209,541 14,122 2,144 15.18
D07 Ropes; cables other than electric 8,943 200 25 12.5
D10 Indexing scheme associated with sublasses of section d, relating to textiles 15,075 1,190 182 15.29
D21 Paper-making; production of cellulose 51,379 7,802 2,098 26.89
D99 Subject matter not otherwise provided for in section d 8 1 1 100
E01 Construction of roads, railways, or bridges 323,633 34,604 6,961 20.12
E02 Hydraulic engineering; foundations; soil shifting 334,875 67,666 10,521 15.55
E03 Water supply; sewerage 193,710 116,767 19,523 16.72
E04 Building 798,420 168,314 65,037 38.64
E05 Locks; keys; window or door fittings; safes 230,626 7,440 1,394 18.74
E06 Doors, windows, shutters, or roller blinds, in general; ladders 205,454 21,416 9,269 43.28
@@
E21 Earth drilling; mining 343,751 29,455 3,886 13.19
E99 Subject matter not otherwise provided for in section e 10 0 0
F01 Machines or engines in general; engine plants in general; steam engines 194,029 91,170 9,797 10.75
F02 Internal-combustion engines; hot-gas or combustion-product engine plants 222,379 119,569 9,673 8.09
F03 Machines or engines for liquids; wind, spring, weight, or miscellaneous motors 102,508 88,037 18,697 21.24
F04 Engines or pumps 317,489 27,271 3,936 14.43
F05 Indexing schemes relating to engines or pumps in various subclasses of classes f01-f04 61,557 34,853 2,594 7.44
F15 Fluid-pressure actuators; hydraulics or pneumatics in general 86,427 4,544 487 10.72
F16 Engineering elements or units; general measures for effecting and maintaining effective functioning of machines 1,413,612 105,839 14,431 13.63
F17 Storing or distributing gases or liquids 71,948 11,868 1,365 11.5
F21 Lighting 482,334 84,536 29,415 34.8
F22 Steam generation 41,347 14,985 4,945 33
F23 Combustion apparatus; combustion processes 157,204 81,676 18,076 22.13
F24 Heating; ranges; ventilating 551,446 198,789 78,356 39.42
F25 Refrigeration or cooling; combined heating and refrigeration systems; heat pump systems 210,456 48,985 12,638 25.8
F26 Drying 201,684 27,834 6,931 24.9
F27 Furnaces; kilns; ovens; retorts 103,060 32,363 8,476 26.19
F28 Heat exchange in general 143,827 39,344 9,540 24.25
F41 Weapons 52,771 1,087 85 7.82
F42 Ammunition; blasting 34,591 1,260 90 7.14
@@
G01 Measuring; testing 2,344,419 171,633 19,988 11.65
G02 Optics 482,028 22,451 2,249 10.02
G03 Photography; cinematography; analogous techniques using waves other than optical waves 246,693 7,362 703 9.55
G04 Horology 34,822 1,531 245 16
G05 Controlling; regulating 367,905 69,208 10,096 14.59
G06 Computing; calculating; counting 1,982,972 187,012 18,944 10.13
G07 Checking-devices 190,290 10,223 2,013 19.69
G08 Signalling 341,630 36,238 6,512 17.97
G09 Educating; cryptography; display; advertising; seals 464,773 27,659 4,892 17.69
G10 Musical instruments; acoustics 120,484 6,012 887 14.75
G11 Information storage 112,154 4,801 153 3.19
G12 Instrument details 4,563 179 20 11.17
G16 Information and communication technology [ict] specially adapted for specific application fields 84,890 8,583 716 8.34
G21 Nuclear physics; nuclear engineering 40,330 22,713 1,258 5.54
@@
H01 Electric elements 2,001,196 625,938 101,705 16.25
H02 Generation, conversion, or distribution of electric power 1,234,173 436,760 155,010 35.49
H03 Electronic circuitry 175,540 11,015 506 4.59
H04 Electric communication technique 1,725,419 126,469 12,398 9.8
H05 Electric techniques not otherwise provided for 656,099 118,791 15,953 13.43
H10 Semiconductor devices; electric solid-state devices not otherwise provided for 131,510 44,066 6,206 14.08
H99 Subject matter not otherwise provided for in section H 22 3 1 33.33
Y02 Technologies or applications for mitigation or adaptation against climate change 2,439,032 2,439,032 531,075 21.77
Y04 Information or communication technologies having an impact on other technology areas 55,239 55,239 9,858 17.85
Y10 Technical subjects covered by former uspc 172,157 29,600 5,189 17.53
@@
A01 Agriculture;forestry; animal husbandry; hunting;trapping; fishing 1.51 1.43
@@
A22 Butchering; meat treatment; processing poultry or fish 0.36 0.21
A23 Foods or foodstuffs; treatment thereof, not covered by other classes 0.58 0.47
A24 Tobacco; cigars; cigarettes; smokers’ requisites 0.27 0.16
@@
A45 Hand or travelling articles 0.17 0.23
A46 Brushware 0.08 0.06
@@
A61 Medical or veterinary science; hygiene 0.28 0.17
A62 Life-saving; fire-fighting 0.59 0.49
@@
A99 Subject matter not otherwise provided for in section a 0.81 0.46
B01 Physical or chemical processes or apparatus in general 2.64 2.37
B02 Crushing, pulverising, or disintegrating; preparatory treatment of grain for milling 0.89 0.83
B03 Separation of solid materials using liquids or using pneumatic tables or jigs 2.36 2.2
B04 Centrifugal apparatus or machines for carrying-out physical or chemical processes 0.86 0.55
@@
B07 Separating solids from solids; sorting 0.53 0.4
B08 Cleaning 0.91 0.9
B09 Disposal of solid waste; reclamation of contaminated soil 5.47 6.81
@@
B22 Casting; powder metallurgy 0.68 0.22
@@
B26 Hand cutting tools; cutting; severing 0.28 0.18
B27 Working or preserving wood or similar material; nailing or stapling machines in general 0.26 0.17
B28 Working cement, clay, or stone 0.42 0.32
B29 Working of plastics; working of substances in a plastic state in general 0.54 0.46
B30 Presses 0.92 0.82
B31 Making paper articles; working paper 0.18 0.12
B32 Layered products 0.6 0.85
B33 Additive manufacturing technology 1.4 0.29
B41 Printing; lining machines; typewriters; stamps 0.2 0.15
B42 Bookbinding; albums; files; special printed matter 0.1 0.03
@@
B44 Decorative arts 0.2 0.17
B60 Vehicles in general 1.11 1.43
B61 Railways 0.55 0.32
B62 Land vehicles for travelling otherwise than on rails 0.38 0.31
B63 Ships or other waterborne vessels; related equipment 1.39 1.29
B64 Aircraft; aviation; cosmonautics 0.88 0.56
@@
B65 Conveying; packing; storing; handling thin or filamentary material 0.49 0.34
B66 Hoisting; lifting; hauling 0.22 0.11
@@
B81 Microstructural devices or systems, e.g. micromechanical devices 0.29 0.08
B82 Nanotechnology 1.51 1.2
B99 Subject matter not otherwise provided for in section b 1.12 0.0
C01 Inorganic chemistry 2.3 2.4
C02 Treatment of water, waste water, sewage, or sludge 5.77 10.37
C03 Glass; mineral or slag wool 1.05 0.72
C04 Cements; concrete; artificial stone; ceramics; refractories 1.72 1.96
C05 Fertilisers; manufacture thereof 3.92 4.69
C06 Explosive; matches 0.31 0.15
C07 Organic chemistry 1.42 0.95
C08 Organic macromolecular compounds; their preparation or chemical working-up 0.8 0.79
C09 Dyes; paints, polishes, natural resins, adhesives; compositions not otherwise provided for 0.96 0.71
C10 Petroleum, gas or coke industries, technical gases containing carbon monoxide; fuels; lubricant; peat 2.34 3.73
C11 Animal or vegetable oils, fats, fatty substances or waxes; fatty acids therefrom; detergents; candles 0.99 1.26
C12 Biochemistry; beer; spirits; wine; vinegar; microbiology; enzymology; mutation or genetic engineering 1.29 1.24
C13 Sugar industry 1.25 3.13
C14 Skins; hides; pelts; leather 0.24 0.18
C21 Metallurgy of iron 1.31 0.86
C22 Metallurgy; ferrous or non-ferrous alloys; treatment of alloys or non-ferrous metals 1.68 1.33

### v2

@@ 05-09 Mining and extraction 59 31 9
10-33 Manifacture 9,771 2,149 366
35 Electricity, gas, steam and air conditioning supply 96 70 20
@@ 36-39 Water, sewerage and waste management 100 69 19
41-43 Construction, civil engineering and special activities 388 118 12
45-47 Wholesale and retail trade 1,520 275 45
49-53 Transport and storage 81 21 4
55-56 Accommodation and food service activities 15 2 0
58-63 Information and communication 450 54 10
64-69 Financial, insurance and real estate activities 391 87 16
69-75 Professional, scientific and technical activities 2,106 691 155
77-79 Administrative and support service activities 88 15 6
80-82 Security and investigation activities 110 29 3
@@ 84 Public administration and defence 4 1 0
85 Education 92 56 25
86-88 Human health and social work activities 88 13 0
90-93 Arts, entertainment and recreation 15 1 0
94-96 Other services activities 32 6 1
@@ B.2. Classification results
A01 Agriculture;forestry; animal husbandry; hunting;trapping; fishing 918,032 246,236 42,903 17.42
A21 Baking; edible doughs 46,668 1,742 363 20.84
A22 Butchering; meat treatment; processing poultry or fish 26,674 1,719 230 13.38
A23 Foods or foodstuffs; treatment thereof, not covered by other classes 340,445 35,337 5,170 14.63
A24 Tobacco; cigars; cigarettes; smokers’ requisites 67,262 3,230 360 11.15
A41 Wearing apparel 150,247 2,535 477 18.82
A42 Headwear 25,378 604 132 21.85
A43 Footwear 74,228 1,211 134 11.07
A44 Haberdashery; jewellery 48,613 864 72 8.33
A45 Hand or travelling articles 172,141 5,337 1,485 27.82
A46 Brushware 32,601 479 79 16.49
A47 Furniture;domestic articles or appliances; coffee mills; spice mills;suction cleaners in general 962,723 34,023 5,318 15.63
A61 Medical or veterinary science; hygiene 1,953,512 99,157 11,076 11.17
A62 Life-saving; fire-fighting 116,193 12,320 1,783 14.47
A63 Sports; games; amusements 376,930 11,391 740 6.5
@@
B01 Physical or chemical processes or apparatus in general 1,332,667 610,307 120,771 19.79
B02 Crushing, pulverising, or disintegrating; preparatory treatment of grain for milling 228,879 36,527 7,753 21.23
B03 Separation of solid materials using liquids or using pneumatic tables or jigs 68,915 29,006 5,558 19.16
B04 Centrifugal apparatus or machines for carrying-out physical or chemical processes 28,461 4,358 624 14.32
B05 Spraying or atomising in general; applying liquids or other fluent materials to surfaces, in general 285,289 31,867 4,378 13.74
B06 Generating or transmitting mechanical vibrations in general 10,705 398 37 9.3
B07 Separating solids from solids; sorting 198,516 18,848 3,138 16.65
B08 Cleaning 437,757 71,364 16,123 22.59
B09 Disposal of solid waste; reclamation of contaminated soil 57,356 55,809 16,145 28.93
B21 Mechanical metal-working without essentially removing material; punching metal 426,389 16,896 1,196 7.08
B22 Casting; powder metallurgy 191,829 23,485 1,319 5.62
B23 Machine tools; metal-working not otherwise provided for 934,950 47,290 3,723 7.87
B24 Grinding; polishing 272,113 14,671 1,408 9.6
B25 Hand tools; portable power-driven tools; tool handles 435,173 15,788 1,096 6.94
B26 Hand cutting tools; cutting; severing 198,229 10,062 1,274 12.66
B27 Working or preserving wood or similar material; nailing or stapling machines in general 83,075 3,921 522 13.31
B28 Working cement, clay, or stone 171,793 12,839 1,735 13.51
B29 Working of plastics; working of substances in a plastic state in general 541,380 52,211 8,806 16.87
B30 Presses 67,870 11,163 2,079 18.62
B31 Making paper articles; working paper 54,618 1,712 230 13.43
B32 Layered products 271,991 29,462 4,588 15.57
B33 Additive manufacturing technology 58,020 14,563 436 2.99
B41 Printing; lining machines; typewriters; stamps 212,727 7,488 1,208 16.13
B42 Bookbinding; albums; files; special printed matter 47,299 870 61 7.01
B43 Writing or drawing implements; bureau accessories 69,639 615 92 14.96
B44 Decorative arts 52,726 1,840 312 16.96
B60 Vehicles in general 952,725 189,225 53,725 28.39
B61 Railways 93,671 9,141 1,076 11.77
B62 Land vehicles for travelling otherwise than on rails 396,508 27,174 5,156 18.97
B63 Ships or other waterborne vessels; related equipment 129,667 32,265 5,760 17.85
@@
B64 Aircraft; aviation; cosmonautics 151,284 23,935 3,208 13.4
B65 Conveying; packing; storing; handling thin or filamentary material 1,509,862 135,778 17,108 12.6
B66 Hoisting; lifting; hauling 340,304 13,225 1,425 10.78
B67 Opening or closing bottles, jars or similar containers; liquid handling 65,340 2,486 206 8.29
B68 Saddlery; upholstery 4,397 82 7 8.54
B81 Microstructural devices or systems, e.g. micromechanical devices 18,449 946 39 4.12
B82 Nanotechnology 71,531 19,314 3,051 15.8
B99 Subject matter not otherwise provided for in section b 5 1 1 100
C01 Inorganic chemistry 186,106 76,276 16,567 21.72
C02 Treatment of water, waste water, sewage, or sludge 424,569 424,558 119,694 28.19
C03 Glass; mineral or slag wool 84,510 15,840 1,973 12.46
C04 Cements; concrete; artificial stone; ceramics; refractories 123,836 37,926 7,804 20.58
C05 Fertilisers; manufacture thereof 43,954 30,734 7,160 23.3
C06 Explosive; matches 7,240 405 40 9.88
C07 Organic chemistry 394,040 99,820 11,371 11.39
C08 Organic macromolecular compounds; their preparation or chemical working-up 383,625 54,682 10,154 18.57
C09 Dyes; paints, polishes, natural resins, adhesives; compositions not otherwise provided for 288,443 49,330 6,766 13.72
C10 Petroleum, gas or coke industries, technical gases containing carbon monoxide; fuels; lubricant; peat 121,248 50,506 17,055 33.77
C11 Animal or vegetable oils, fats, fatty substances or waxes; fatty acids therefrom; detergents; candles 49,336 8,744 2,341 26.77
C12 Biochemistry; beer; spirits; wine; vinegar; microbiology; enzymology; mutation or genetic engineering 369,128 85,351 10,731 12.57
C13 Sugar industry 4,212 942 467 49.58
C14 Skins; hides; pelts; leather 9,732 422 74 17.54
C21 Metallurgy of iron 120,910 28,231 3,486 12.35
C22 Metallurgy; ferrous or non-ferrous alloys; treatment of alloys or non-ferrous metals 144,029 43,131 7,541 17.48
C23 Coating metallic material; coating material with metallic material; chemical surface treatment 159,300 21,953 2,770 12.62
C25 Electrolytic or electrophoretic processes; apparatus therefor 92,938 25,917 4,617 17.81
C30 Crystal growth 31,669 12,000 1,213 10.11
C40 Combinatorial chemistry; libraries, e.g. chemical libraries, dna libraries 4,883 570 20 3.51
@@
D01 Natural or man-made threads or fibres; spinning 84,326 6,582 1,126 17.11
D02 Yarns; mechanical finishing of yarns or ropes; warping or beaming 26,964 1,039 133 12.8
D03 Weaving 40,326 1,176 131 11.14
D04 Braiding; lace-making; knitting; trimmings; non-woven fabrics 58,316 3,605 393 10.9
D05 Sewing; embroidering; tufting 42,533 548 22 4.01
D06 Treatment of textiles or the like; laundering; flexible materials not otherwise provided for 209,554 14,122 2,191 15.51
D07 Ropes; cables other than electric 8,945 200 16 8
D10 Indexing scheme associated with sublasses of section d, relating to textiles 15,129 1,199 164 13.68
D21 Paper-making; production of cellulose 51,382 7,804 2,050 26.27
D99 Subject matter not otherwise provided for in section d 9 1 1 100
E01 Construction of roads, railways, or bridges 323,643 34,608 6,812 19.68
E02 Hydraulic engineering; foundations; soil shifting 334,899 67,676 9,971 14.73
E03 Water supply; sewerage 193,719 116,768 18,110 15.51
E04 Building 798,477 168,343 29,548 17.55
E05 Locks; keys; window or door fittings; safes 230,644 7,442 966 12.98
E06 Doors, windows, shutters, or roller blinds, in general; ladders 205,466 21,419 4,070 19
E21 Earth drilling; mining 343,772 29,461 2,757 9.36
@@
E99 Subject matter not otherwise provided for in section E 10 0 0 0
F01 Machines or engines in general; engine plants in general; steam engines 194,068 91,189 10,496 11.51
F02 Internal-combustion engines; hot-gas or combustion-product engine plants 222,402 119,576 10,353 8.66
F03 Machines or engines for liquids; wind, spring, weight, or miscellaneous motors 102,544 88,059 18,877 21.44
F04 Engines or pumps 317,546 27,282 4,048 14.84
F05 Indexing schemes relating to engines or pumps in various subclasses of classes f01-f04 61,699 35,014 2,689 7.68
F15 Fluid-pressure actuators; hydraulics or pneumatics in general 86,451 4,551 536 11.78
F16 Engineering elements or units; general measures for effecting and maintaining effective functioning of machines 1,413,801 105,884 13,355 12.61
F17 Storing or distributing gases or liquids 71,961 11,870 931 7.84
F21 Lighting 482,428 84,562 29,276 34.62
F22 Steam generation 41,352 14,987 5,195 34.66
F23 Combustion apparatus; combustion processes 157,212 81,683 20,375 24.94
F24 Heating; ranges; ventilating 551,508 198,801 76,181 38.32
F25 Refrigeration or cooling; combined heating and refrigeration systems; heat pump systems 210,520 49,008 12,255 25.01
F26 Drying 201,694 27,839 7,292 26.19
F27 Furnaces; kilns; ovens; retorts 103,068 32,364 9,155 28.29
F28 Heat exchange in general 143,847 39,355 9,325 23.69
F41 Weapons 52,780 1,088 65 5.97
F42 Ammunition; blasting 34,595 1,260 74 5.87
@@
G01 Measuring; testing 2,344,888 171,704 16,917 9.85
G02 Optics 482,153 22,468 1,966 8.75
G03 Photography; cinematography; analogous techniques using waves other than optical waves 246,789 7,373 541 7.34
G04 Horology 34,830 1,533 249 16.24
G05 Controlling; regulating 368,198 69,261 9,241 13.34
G06 Computing; calculating; counting 1,983,168 187,084 14,218 7.6
G07 Checking-devices 190,431 10,242 1,897 18.52
G08 Signalling 341,739 36,256 6,471 17.85
G09 Educating; cryptography; display; advertising; seals 464,858 27,671 4,451 16.09
G10 Musical instruments; acoustics 120,526 6,015 516 8.58
G11 Information storage 112,199 4,806 122 2.54
G12 Instrument details 4,565 179 17 9.5
G16 Information and communication technology [ict] specially adapted for specific application fields 85,329 8,602 451 5.24
G21 Nuclear physics; nuclear engineering 40,347 22,718 755 3.32
@@
H01 Electric elements 1,961,055 610,482 84,510 13.84
H02 Generation, conversion, or distribution of electric power 1,234,321 436,827 156,797 35.89
H03 Electronic circuitry 175,601 11,017 441 4
H04 Electric communication technique 1,726,290 126,570 11,399 9.01
H05 Electric techniques not otherwise provided for 656,296 118,823 13,972 11.76
H10 Semiconductor devices; electric solid-state devices not otherwise provided for 301,283 120,771 27,454 22.73
H99 Subject matter not otherwise provided for in section H 24 3 1 33.33
Y02 Technologies or applications for mitigation or adaptation against climate change 2,442,727 2,439,432 489,298 20.06
Y04 Information or communication technologies having an impact on other technology areas 55,317 55,249 9,227 16.7
Y10 Technical subjects covered by former uspc 172,307 29,631 4,990 16.84
@@
A01 Agriculture;forestry; animal husbandry; hunting;trapping; fishing 1.5 1.36
@@
A22 Butchering; meat treatment; processing poultry or fish 0.36 0.25
A23 Foods or foodstuffs; treatment thereof, not covered by other classes 0.58 0.44
A24 Tobacco; cigars; cigarettes; smokers’ requisites 0.27 0.15
@@
A45 Hand or travelling articles 0.17 0.25
A46 Brushware 0.08 0.07
@@
A61 Medical or veterinary science; hygiene 0.27 0.16
A62 Life-saving; fire-fighting 0.59 0.44
@@
A99 Subject matter not otherwise provided for in section a 0.8 0.5
B01 Physical or chemical processes or apparatus in general 2.63 2.73
B02 Crushing, pulverising, or disintegrating; preparatory treatment of grain for milling 0.89 0.98
B03 Separation of solid materials using liquids or using pneumatic tables or jigs 2.35 2.33
B04 Centrifugal apparatus or machines for carrying-out physical or chemical processes 0.85 0.63
@@
B07 Separating solids from solids; sorting 0.53 0.46
B08 Cleaning 0.91 1.06
B09 Disposal of solid waste; reclamation of contaminated soil 5.45 8.2
@@
B22 Casting; powder metallurgy 0.68 0.2
@@
B26 Hand cutting tools; cutting; severing 0.28 0.19
B27 Working or preserving wood or similar material; nailing or stapling machines in general 0.26 0.18
B28 Working cement, clay, or stone 0.42 0.29
B29 Working of plastics; working of substances in a plastic state in general 0.54 0.47
B30 Presses 0.92 0.88
B31 Making paper articles; working paper 0.17 0.12
B32 Layered products 0.6 0.49
B33 Additive manufacturing technology 1.4 0.22
B41 Printing; lining machines; typewriters; stamps 0.2 0.16
B42 Bookbinding; albums; files; special printed matter 0.1 0.04
@@
B44 Decorative arts 0.19 0.17
B60 Vehicles in general 1.11 1.65
B61 Railways 0.54 0.33
B62 Land vehicles for travelling otherwise than on rails 0.38 0.37
B63 Ships or other waterborne vessels; related equipment 1.39 1.28
B64 Aircraft; aviation; cosmonautics 0.88 0.61
@@
B65 Conveying; packing; storing; handling thin or filamentary material 0.49 0.32
B66 Hoisting; lifting; hauling 0.22 0.12
@@
B81 Microstructural devices or systems, e.g. micromechanical devices 0.29 0.06
B82 Nanotechnology 1.51 1.23
B99 Subject matter not otherwise provided for in section b 1.12 5.78
C01 Inorganic chemistry 2.29 2.59
C02 Treatment of water, waste water, sewage, or sludge 5.76 8.64
C03 Glass; mineral or slag wool 1.05 0.67
C04 Cements; concrete; artificial stone; ceramics; refractories 1.71 1.82
C05 Fertilisers; manufacture thereof 3.91 4.72
C06 Explosive; matches 0.31 0.16
C07 Organic chemistry 1.42 0.83
C08 Organic macromolecular compounds; their preparation or chemical working-up 0.79 0.76
C09 Dyes; paints, polishes, natural resins, adhesives; compositions not otherwise provided for 0.95 0.68
C10 Petroleum, gas or coke industries, technical gases containing carbon monoxide; fuels; lubricant; peat 2.33 4.09
C11 Animal or vegetable oils, fats, fatty substances or waxes; fatty acids therefrom; detergents; candles 0.99 1.37
C12 Biochemistry; beer; spirits; wine; vinegar; microbiology; enzymology; mutation or genetic engineering 1.29 0.84
C13 Sugar industry 1.25 3.2
C14 Skins; hides; pelts; leather 0.24 0.22
C21 Metallurgy of iron 1.3 0.83
C22 Metallurgy; ferrous or non-ferrous alloys; treatment of alloys or non-ferrous metals 1.67 1.51